# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@
Antonin Boisneault    Marcella Bonazzoli    Xavier Claeys    Pierre Marchand 1122
@@ 1 Introduction and definition of the problem
When solving the Helmholtz equation in complex heterogeneous media, it is of interest to decompose the domain according to the variation of the wavenumber. Local problems in homogeneous subdomains can be reformulated as equations set on their boundary (which is particularly useful for unbounded subdomains), and the global problem is solved with Finite Element Method - Boundary Element Method (FEM-BEM) coupling techniques BielakMacCamy1983EIP; Costabel1987SMC; JohnsonNedelec1980CBI. Recently, a substructured formulation, called Generalized Optimized Schwarz Method (GOSM), has been designed for bounded domains, with weakly imposed boundary conditions, see Claeys2023NOS, and extended in ourPaper to unbounded domains and interface conditions arising from several FEM-BEM coupling techniques. Unfortunately, FEM-BEM variational formulations can be ill-posed for certain wavenumbers called spurious resonances SchulzHiptmair2022SRC. In this paper, we focus on the Johnson-Nédélec JohnsonNedelec1980CBI and Costabel Costabel1987SMC couplings and show that the associated GOSMs suffer from that issue. For both couplings, we give an explicit expression of the kernel of the local operator associated with the interface between the FEM and BEM subdomains. This kernel and the one of the corresponding classical FEM-BEM couplings are simultaneously non-trivial.
@@
In FEM-BEM coupling, the FEM is used for the (possibly) heterogeneous subdomain $\Omega_{\text{F}}$, the BEM is applied for the homogeneous domain $\Omega_{\text{B}}$, and transmission conditions are imposed at the interface $\Sigma\coloneqq\partial\Omega_{\text{F}}\cap\partial\Omega_{\text{B}}$. Our FEM-BEM formulation, based on the Generalized Optimized Schwarz Method (GOSM) ourPaper; Claeys2023NOS, is set on $\Sigma$ and writes:
@@
where $\operatorname{\Pi}$ is a (possibly) non-local exchange operator, $\operatorname{S}\coloneqq\operatorname{diag}(\operatorname{S}_{\text{B}},\operatorname{S}_{\text{F}})$ is a block-diagonal scattering operator and $q_{\text{B}},q_{\text{F}}\in H^{-1/2}(\Sigma)$ are outgoing impedance traces shared between $\Omega_{\text{B}}$ and $\Omega_{\text{F}}$. More details about (2) can be found in (Claeys2023NOS, Propos. 8.1). To derive the substructured formulation (2), the starting point is to write Problem (1) in variational form
@@
$\int_{\Omega_{\text{F}}}(\nabla u\cdot\nabla v-\kappa^{2}uv)\mathop{}\!\mathrm{d}\bm{\bm{x}}+\langle\operatorname{A}_{\Sigma}(u|_{\Sigma},p),(v|_{\Sigma},q)\rangle=l_{\Omega_{\text{F}}}(v),$
where the explicit expression of $\operatorname{A}_{\Sigma}\colon H^{{1}/{2}}(\Sigma)\times H^{-{1}/{2}}(\Sigma)\to H^{-{1}/{2}}(\Sigma)\times H^{{1}/{2}}(\Sigma)$ depends on the choice of the FEM-BEM coupling and involves Boundary Integral Operators (BIOs) (see Section 2). The linear form $l_{\Omega_{\text{F}}}$ accounts for the contributions of the source term $f$ and the boundary condition on $\Omega_{\text{O}}$. Here, the canonical duality pairing between a Banach space $H$ and its topological dual $H^{*}$ is denoted $\langle\cdot,\cdot\rangle\colon H^{*}\times H\to\mathbb{C}$ and defined by $\langle\varphi,v\rangle\coloneq\varphi(v)$. We emphasize that the duality pairings we consider do not involve any complex conjugation.
Next, we introduce a boundary operator $\operatorname{B}_{\Sigma}\colon H^{{1}/{2}}(\Sigma)\times H^{-{1}/{2}}(\Sigma)\to H^{{1}/{2}}(\Sigma)$, defined by $\operatorname{B}_{\Sigma}(\phi,p)\coloneqq\phi$, and also consider a transmission (or impedance) operator $\operatorname{T}_{\Sigma}\colon\!H^{{1}/{2}}(\Sigma)\to H^{-{1}/{2}}(\Sigma)$, satisfying $\langle\operatorname{T}_{\Sigma}(\phi),\overline{\phi}\rangle>0\;\;\forall\phi\in H^{1/2}(\Sigma)\setminus\{0\}$, $\operatorname{T}_{\Sigma}^{*}=\operatorname{T}_{\Sigma}$, and $\operatorname{Re}(\operatorname{T}_{\Sigma})>0$. The scattering operator $\operatorname{S}_{\text{B}}$ involved in the GOSM is given by $\operatorname{Id}+2\,\imath\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma}(\operatorname{A}_{\Sigma}-\imath\operatorname{B}_{\Sigma}^{*}\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma})^{-1}\operatorname{B}_{\Sigma}^{*}$. We aim to study the kernel of $\operatorname{A}_{\Sigma}-\imath\operatorname{B}_{\Sigma}^{*}\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma}$, which is crucial to establishing the well-posedness of the GOSM.
@@ 3 Numerical illustration
The spurious resonances for the considered geometry are the zeros of the first kind Bessel functions. We assume that $\kappa$ is not a spurious resonance, so $\operatorname{A}_{\Sigma}-\imath\operatorname{B}^{*}_{\Sigma}\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma}$ is invertible. To illustrate the sensitivity or the robustness to spurious resonances of the GOSM for the Johnson-Nédélec and Costabel couplings, we study how the relative error of several quantities of interest evolves with respect to $\kappa$. We focus on the wavenumber range $\kappa\in[4.28,4.42]$, inside which only $\kappa_{1}\approx 4.32685$ and $\kappa_{2}\approx 4.38575$ are spurious resonances. All the experiments have been led with a fixed mesh generated for $\kappa=10$ and $20$ points per wavelength.
@@
We now turn to the evolution of the relative errors of $\gamma_{D}(u)$ and $\gamma_{N}(u)$, with respect to $\kappa$. We emphasize that $\gamma_{D}(u)$ and $\gamma_{N}(u)$ are reconstructed from the solution of the GOSM, namely $\left(\gamma_{D}(u),\gamma_{N}(u)\right)\coloneqq(\operatorname{A}_{\Sigma}-\imath\operatorname{B}_{\Sigma}^{*}\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma})^{-1}\operatorname{B}_{\Sigma}^{*}q_{\text{B}}=v_{q_{\text{B}}}.$ In Figure 3, when $\kappa$ is close to a spurious resonance we observe peaks only in the Neumann relative error curve for the Costabel coupling, while there are peaks in both Dirichlet and Neumann curves for the Johnson-Nédélec coupling. The results confirm what we said previously about the deterioration near a spurious resonance of the components of $v_{q}$ for a given vector $q$. It is interesting to note that the reconstructed solution $(\gamma_{D}(u),\gamma_{N}(u))$ is more and more polluted when $\kappa$ becomes closer to a spurious resonance, even though $\operatorname{A}_{\Sigma}-\imath\operatorname{B}^{*}_{\Sigma}\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma}$ is invertible. We highlight that the relative errors go from $1\%$ to more than $100\%$, but are not comparable: around $\kappa_{1}$ the Dirichlet relative error is close to $1$, while the Neumann relative error is close to $0.4$. These two errors are too large for $q_{\text{B}}$ and the reconstructed solution to be considered as good quality approximations.
@@
Finally, we pinpoint that relative errors peaks arise in the same way when classical FEM-BEM couplings (3) are considered. Indeed, for a geometry with two subdomains without an obstacle, it has been proven that only the Neumann component of the Costabel coupling can be polluted, see for instance (SchulzHiptmair2022SRC, Theorem 1), while for the Johnson-Nédélec coupling both volume and Neumann components can be polluted.

### v2

@@
[…]
1122
@@ 1 Introduction and definition of the problem
When solving the Helmholtz equation in complex heterogeneous media, it is of interest to decompose the domain according to the variation of the wavenumber. Local problems in homogeneous subdomains can be reformulated as equations set on their boundary (which is particularly useful for unbounded subdomains), and the global problem is solved with Finite Element Method - Boundary Element Method (FEM-BEM) coupling techniques BielakMacCamy1983EIP; Costabel1987SMC; JohnsonNedelec1980CBI. Recently, a substructured formulation, called Generalized Optimized Schwarz Method (GOSM), has been designed for bounded domains, with weakly imposed boundary conditions, see Claeys2023NOS, and extended in ourPaper to unbounded domains and interface conditions arising from several FEM-BEM coupling techniques. Unfortunately, FEM-BEM variational formulations can be ill-posed for certain wavenumbers called spurious resonances SchulzHiptmair2022SRC. The combined FEM-BEM formulations from HiptmairMeury2006SFB eliminate the spurious resonances, but rely on a compact regularizing operator and their numerical implementation might be complicated. In this paper, we choose to focus on the classical Johnson-Nédélec JohnsonNedelec1980CBI and Costabel Costabel1987SMC couplings and show that also the associated GOSMs suffer from the same spurious resonances. For both couplings, we give an explicit expression of the kernel of the local operator associated with the interface between the FEM and BEM subdomains. This kernel and the one of the corresponding classical FEM-BEM couplings are simultaneously non-trivial.
@@
In FEM-BEM coupling, the FEM is used for the (possibly) heterogeneous subdomain $\Omega_{\text{F}}$, the BEM is applied for the homogeneous domain $\Omega_{\text{B}}$, and transmission conditions are imposed at the interface $\Sigma\coloneqq\partial\Omega_{\text{F}}\cap\partial\Omega_{\text{B}}$. Our FEM-BEM formulation, based on the Generalized Optimized Schwarz Method (GOSM) ourPaper; Claeys2023NOS, is set on $\Sigma$ and reads:
@@
where $\operatorname{\Pi}$ is a (possibly) non-local exchange operator, $\operatorname{S}\coloneqq\operatorname{diag}(\operatorname{S}_{\text{B}},\operatorname{S}_{\text{F}})$ is a block-diagonal scattering operator and $q_{\text{B}},q_{\text{F}}\in H^{-1/2}(\Sigma)$ are outgoing impedance traces shared between $\Omega_{\text{B}}$ and $\Omega_{\text{F}}$. More details about (2) can be found in (Claeys2023NOS, Prop. 8.1). To derive the substructured formulation (2), the starting point is to write Problem (1) in variational form
@@
$\int_{\Omega_{\text{F}}}(\nabla u\cdot\nabla v-\kappa^{2}uv)\mathop{}\!\mathrm{d}\bm{\bm{x}}+\langle\operatorname{A}_{\Sigma}(u|_{\Sigma},p),(v|_{\Sigma},q)\rangle=\ell(v),$
where the explicit expression of $\operatorname{A}_{\Sigma}\colon H^{{1}/{2}}(\Sigma)\times H^{-{1}/{2}}(\Sigma)\to H^{-{1}/{2}}(\Sigma)\times H^{{1}/{2}}(\Sigma)$ depends on the choice of the FEM-BEM coupling and involves Boundary Integral Operators (BIOs) (see Section 2). The linear form $\ell$ accounts for the contributions of the source term $f$ and the boundary condition on $\Omega_{\text{O}}$. For instance, $\ell(v)=\int_{\Omega_{\text{F}}}fv\mathop{}\!\mathrm{d}\bm{\bm{x}}$ for homogeneous Neumann boundary condition, or if $\Omega_{\text{O}}=\emptyset$. Here, the canonical duality pairing between a Banach space $H$ and its topological dual $H^{*}$ is denoted $\langle\cdot,\cdot\rangle\colon H^{*}\times H\to\mathbb{C}$ and defined by $\langle\varphi,v\rangle\coloneq\varphi(v)$. We emphasize that the duality pairings we consider do not involve any complex conjugation.
Next, we introduce a boundary operator $\operatorname{B}_{\Sigma}\colon H^{{1}/{2}}(\Sigma)\times H^{-{1}/{2}}(\Sigma)\to H^{{1}/{2}}(\Sigma)$, defined by $\operatorname{B}_{\Sigma}(\phi,p)\coloneqq\phi$, and also consider a transmission (or impedance) operator $\operatorname{T}_{\Sigma}\colon\!H^{{1}/{2}}(\Sigma)\to H^{-{1}/{2}}(\Sigma)$, satisfying $\langle\operatorname{T}_{\Sigma}(\phi),\overline{\phi}\rangle>0\;\;\forall\phi\in H^{1/2}(\Sigma)\setminus\{0\}$, $\operatorname{T}_{\Sigma}^{*}=\operatorname{T}_{\Sigma}$, and $\operatorname{Re}(\operatorname{T}_{\Sigma})>0$. The scattering operator $\operatorname{S}_{\text{B}}$ involved in the GOSM is given by $\operatorname{Id}+2\,\imath\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma}(\operatorname{A}_{\Sigma}-\imath\operatorname{B}_{\Sigma}^{*}\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma})^{-1}\operatorname{B}_{\Sigma}^{*}$, see (Claeys2023NOS, Prop. 7.2). We aim to study the kernel of $\operatorname{A}_{\Sigma}-\imath\operatorname{B}_{\Sigma}^{*}\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma}$, which is crucial to establishing the well-posedness of the GOSM.
@@ 3 Numerical illustration
The spurious resonances for the considered geometry are the zeros of $J_{\nu}$ divided by $2$ ($\Omega_{\text{F}}$ is of radius $2$), see e.g. (ChandlerWildeGrahamEtAl2012NAB, Th. 2.25). We assume that $\kappa$ is not a spurious resonance, so $\operatorname{A}_{\Sigma}-\imath\operatorname{B}^{*}_{\Sigma}\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma}$ is invertible. To illustrate the sensitivity or the robustness to spurious resonances of the GOSM for the Johnson-Nédélec and Costabel couplings, we study how the relative error of several quantities of interest evolves with respect to $\kappa$. We focus on the wavenumber range $\kappa\in[4.28,4.42]$, inside which only $\kappa_{1}\approx 4.32685$ and $\kappa_{2}\approx 4.38575$ are spurious resonances. All the experiments have been led with a fixed mesh generated for $\kappa=10$ and $20$ points per wavelength.
@@
We now turn to the evolution of the relative errors of $\gamma_{D}(u)$ and $\gamma_{N}(u)$, with respect to $\kappa$. We emphasize that $\gamma_{D}(u)$ and $\gamma_{N}(u)$ are reconstructed from the solution of the GOSM, namely $\left(\gamma_{D}(u),\gamma_{N}(u)\right)\coloneqq(\operatorname{A}_{\Sigma}-\imath\operatorname{B}_{\Sigma}^{*}\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma})^{-1}\operatorname{B}_{\Sigma}^{*}q_{\text{B}}=v_{q_{\text{B}}}.$ In Figure 3, when $\kappa$ is close to a spurious resonance we observe peaks only in the Neumann relative error curve for the Costabel coupling, while there are peaks in both Dirichlet and Neumann curves for the Johnson-Nédélec coupling. The results confirm what we said previously about the deterioration near a spurious resonance of the components of $v_{q}$ for a given vector $q$. It is interesting to note that the reconstructed solution $(\gamma_{D}(u),\gamma_{N}(u))$ is more and more spoiled when $\kappa$ becomes closer to a spurious resonance, even though $\operatorname{A}_{\Sigma}-\imath\operatorname{B}^{*}_{\Sigma}\operatorname{T}_{\Sigma}\operatorname{B}_{\Sigma}$ is invertible. We highlight that the relative errors go from $1\%$ to more than $100\%$, but are not comparable: around $\kappa_{1}$ the Dirichlet relative error is close to $1$, while the Neumann relative error is close to $0.4$. These two errors are too large for $q_{\text{B}}$ and the reconstructed solution to be considered as good quality approximations.
@@
Finally, we pinpoint that relative errors peaks arise in the same way when classical FEM-BEM couplings (3) are considered. Indeed, for a geometry with two subdomains without an obstacle, it has been proven that only the Neumann component of the Costabel coupling can be spoiled, see for instance (SchulzHiptmair2022SRC, Th. 1), while for the Johnson-Nédélec coupling both volume and Neumann components can be spoiled.
[…]
AcknowledgementsThis work is funded by the Inria program “Actions exploratoires” (OptiGPR3D).